OpenStudy (anonymous):

What is the simplified form of x to the seventh power over y to the sixth power all over x-squared over y-cubed?

OpenStudy (anonymous):

\[\frac{ \frac{ x^{7} }{ y^{6} } }{ \frac{ x^{2} }{y^{3} } }\]

OpenStudy (anonymous):